# Aurora-M: The First Open Source Multilingual Language Model Red-teamed according to the U.S. Executive Order

### v2

@@ VI HI VI HI VI HI VI HI VI HI
VinaLlama-7b (Nguyen et al., 2023a) 28.63 18.75 37.39 26.31 27.15 24.12 43.13 39.11 34.07 27.07
@@
Code Evaluation Datasets.
Safety Evaluation Datasets.
For our safety evaluation, we employ the evaluation suite provided by Bianchi et al. (2024) to measure safety across various dimensions. Moreover, we constructed our own 40 English Biden-Harris concerned focused instructions in the categories of privacy, misinformation, harm promotion, malware, CNBR, illegal acts, and cyber attacks. Then we translated these to the other languages, resulting in 280 instructions, which we call the Biden-Harris Redteam Testset. Additionally, we use the DangerousQA dataset (Bhardwaj & Poria, 2023) to measure the Attack Success Rate (ASR) of harmful queries when provided as input to both our base and red-teamed models.
@@ 5.2 Evaluation Results
Tables 5 and 6 illustrate the proficiency of Aurora-M in code generation, demonstrating the possibility of continual pre-training from a code-centric checkpoint on multilingual data. In Table 5, the HumanEval and MBPP evaluation benchmarks assess the model’s ability to generate syntactically and semantically correct code snippets. Aurora-M exhibits competitive performance on the Pass@1 metric, which evaluates the model’s ability to produce a correct answer on the first attempt. In particular, Aurora-M consistently matches or outperforms StarCoderPlus, suggesting a significant improvement in code synthesis capabilities. In Appendix D, we show results on additional code datasets and further analyze the behavior of our system by looking at the relationship between its performance and the number of training tokens across various languages and modalities.
(a) Harmfulness scores of our base model (red) compared to its instruction-tuned version (blue). The lower the better.
Figure 3: Overall safety results.
@@
In Figure 3, we provide the safety results comparing our base model against our Biden-Harris red-teamed model obtained by instruction-tuning the former on the dataset introduced in Section 4. For the Biden-Harris Redteam Testset evaluation, four volunteers reviewed both models’ responses and scored them with -2 if harmful, 1 if not helpful but harmless, and 2 if both helpful and harmless. We term the percentage of the total score per category compared to its maximum possible score as the Continual Alignment Redteam Percentage ("CARP"). We can immediately appreciate the considerably lower harmfulness both on the existing benchmarks and on our own Biden-Harris red-team test set as evident by the CARP scores obtained by our red-teamed Aurora-M. We also note that even though our instruction set is predominantly in English, safety consistently improved not only in our target languages but also in languages we did not specifically focus on, such as German, thus showing strong indications of cross-lingual red-teaming effects. Furthermore, as shown in Appendix D, the Attack Success Rate (ASR) on DangerousQA was also reduced.
@@ 6 Related Work
Expanding Multilingual Language Models. Initially, the development of LLMs has predominantly targeted the English language (Brown et al., 2020), leveraging the extensive corpus of English data available on the Web and the broad applicability of models trained on English text. However, this emphasis has often come at the cost of accommodating the linguistic diversity found across various language demographics (Zhu et al., 2023b; Bang et al., 2023; Zhang et al., 2024). Recognizing this significant limitation (Robinson et al., 2023; Peng et al., 2024), recent research has proposed foundational LLMs equipped with multilingual capabilities (Scao et al., 2023; Chai et al., 2023; Wei et al., 2023; Pires et al., 2019; Lin et al., 2021; Shliazhko et al., 2022), or has explicitly concentrated on addressing the challenges posed by low-resource languages (Üstün et al., 2024; Singh et al., 2024; Gala et al., 2023). To integrate multilingual capabilities into existing LLMs, researchers have proposed a variety of methods to enhance multilingual adaptation. These approaches range from continual pretraining techniques (Ibrahim et al., 2024; Gupta et al., 2023) to initial training on extensive multilingual datasets (Scao et al., 2023; Chai et al., 2023) and then subsequent specialized fine-tuning on a target language (Yang et al., 2023), and even adaptation through instruction tuning (Shaham et al., 2024; Kew et al., 2023; Gala et al., 2024). Critical aspects in multilingual adaptation remain on the availability of high-quality diverse multilingual corpus (Corrêa et al., 2024) and further the scope of vocabulary of the specific language.
Continual Pretraining. The notion of large static datasets becomes impractical when confronted with the dynamic nature of real-time events, evolving facts, and the introduction of new data or concepts within a domain. In such scenarios, continual learning (Ring, 1998; Thrun, 1998; Kirkpatrick et al., 2017; Zenke et al., 2017; Rebuffi et al., 2017; Lopez-Paz & Ranzato, 2017; Yadav et al., 2023a; b; c) becomes imperative for large pretrained models to swiftly adapt to these shifting environments. The drive for continual pretraining (Gururangan et al., 2020) extends beyond the dynamic nature of real-world data; it is also fueled by the prohibitive expense associated with training current foundation models from scratch. As more curated datasets become accessible (e.g. RedPajama (Together, 2023), Dolma (Soldaini et al., 2024), CommonCorpus [^7] https://huggingface.co/blog/Pclanglais/common-corpus, ToolQA (Zhuang et al., 2023)) the idea of incorporating knowledge from these datasets through retraining on the union of all available data sets becomes inherently unfeasible. This also encompasses the integration of new capabilities into foundational models trained on specific data distributions. For example, it involves expanding natural language-based models to include the structured modality of code (e.g. StableCode [^8] https://stability.ai/news/stable-code-2024-llm-code-completion-release). Previous studies on continual pretraining have primarily concentrated on replay, optimizing the learning rate schedule (Ibrahim et al., 2024), preserving general knowledge through soft masking attention heads (Ke et al., 2023), exploring the effects of domain similarity and model capacity on forward and backward transfer (Yıldız et al., 2024), and continual post-training for few-shot adaptation (Ke et al., 2022).
LLM Compliance The extensive utilization of Large Language Models (LLMs) across various applications underscores the necessity for their operation to uphold user privacy, mitigate risks such as misinformation or biased outputs, and ensure transparency regarding their functionality and utilization, all while adhering to local regulations. Independent evaluations and red teaming play crucial roles in assessing these risks. However, conflicts of interest within major AI companies may impede such safety evaluations, underscoring the necessity for a safe harbor for safety research (Longpre et al., 2024). Benchmark datasets facilitating such research efforts include those proposed by Zhang et al. (2023) and Sun et al. (2023). Studies aimed at enhancing the security, safety, and legal compliance of Large Language Models (LLMs) have encompassed various approaches. These include the creation of datasets based on FAIR data principles (Raza et al., 2024), the development of structured LLM auditing mechanisms (Mökander et al., 2023), risk assessment of LLM alignment through personalized feedback (Kirk et al., 2023), structured evaluation of risks associated with LLM deployment (Derczynski et al., 2023), and profiling foundation model transparency (Bommasani et al., 2023; 2024).
@@ 7 Conclusion
We introduce Aurora-M, designed to align with legal standards and enhance accessibility. This model showcases proficiency in multilingual understanding and coding tasks, while prioritizing compliance with the safety guidelines outlined in the Biden-Harris US Executive Order on AI. Moreover, Aurora-M exemplifies the collaborative nature of the open-source community, promoting transparency and accessibility in AI development. By red-teaming the model in accordance with the Biden-Harris Executive Order, it underscores the significance of responsible AI development and ensures alignment with government standards. Striking a balance between safety, utility, and cutting-edge performance, Aurora-M emerges as a valuable tool for researchers and developers. In addition, we present intriguing insights into cross-lingual red-teaming effects and emphasize the importance of the two-stage curriculum-based continual pretraining approach.
Future Work.
Leveraging insights from Aurora-M’s development, we plan to explore continual pretraining of stronger base models using the same two-stage curriculum, prioritizing safety. This applies to both LLMs and Multimodal-LLMs (MLLMs). Furthermore, we aim to train multiple independent domain experts based on Aurora-M, potentially merging them to improve task specialization.
Ethics statement

### v3

@@ VI HI VI HI VI HI VI HI VI HI
ViGPTQA-6b (Nguyen et al., 2023a) - - - - - - 43.26 - - -
VinaLlama-7b (Nguyen et al., 2023b) 28.63 18.75 37.39 26.31 27.15 24.12 43.13 39.11 34.07 27.07
@@
(a) Harmfulness scores of our base model (pink) compared to its instruction-tuned version (blue). The lower the better.
Figure 3: Overall safety results.
@@ 5.2 Evaluation Results
Tables 5 and 6 illustrate the proficiency of Aurora-M in code generation, demonstrating the possibility of continual pre-training from a code-centric checkpoint on multilingual data. In Table 5, the HumanEval and MBPP evaluation benchmarks assess the model’s ability to generate syntactically and semantically correct code snippets. Aurora-M exhibits competitive performance on the Pass@1 metric, which evaluates the model’s ability to produce a correct answer on the first attempt. In particular, Aurora-M consistently matches or outperforms StarCoderPlus, suggesting a significant improvement in code synthesis capabilities. In Appendix E.1, we show results on additional code datasets and further analyze the behavior of our system by looking at the relationship between its performance and the number of training tokens across various languages and modalities.
@@
In Figure 3, we provide the safety results comparing our base model against our Biden-Harris red-teamed model obtained by instruction-tuning the former on the dataset introduced in Section 4. For the Biden-Harris Redteam Testset evaluation, four volunteers reviewed both models’ responses and scored them with -2 if harmful, 1 if not helpful but harmless, and 2 if both helpful and harmless. We term the percentage of the total score per category compared to its maximum possible score as the Continual Alignment Redteam Percentage ("CARP"). We can immediately appreciate the considerably lower harmfulness both on the existing benchmarks and on our own Biden-Harris red-team test set as evident by the CARP scores obtained by our red-teamed Aurora-M. We also note that even though our instruction set is predominantly in English, safety consistently improved not only in our target languages but also in languages we did not specifically focus on, such as German, thus showing strong indications of cross-lingual red-teaming effects. Furthermore, as shown in Appendix E.1, the Attack Success Rate (ASR) on DangerousQA was also reduced.
5.3 Training Analysis
Figure 5 and 6 show the relationship between the number of training tokens and the performance of the various models. This analysis aims to capture these trends for the code generation tasks such as HumanEval and MBPP, as well as for the English, Finnish, Hindi, Japanese, and Vietnamese language evaluations. We refer to Appendix E.2 for detailed discussion.
@@ 6 Related Work
Expanding Multilingual Language Models. Initially, the development of LLMs has predominantly targeted the English language (Brown et al., 2020), leveraging the extensive corpus of English data available on the Web and the broad applicability of models trained on English text. However, this emphasis has often come at the cost of accommodating the linguistic diversity found across various language demographics (Zhu et al., 2023b; Bang et al., 2023; Zhang et al., 2024). Recognizing this significant limitation (Robinson et al., 2023; Peng et al., 2024), recent research has proposed foundational LLMs equipped with multilingual capabilities (Chai et al., 2023; Scao et al., 2023; Wei et al., 2023; Shliazhko et al., 2022), or has explicitly concentrated on addressing the challenges posed by low-resource languages (Üstün et al., 2024; Singh et al., 2024; Gala et al., 2023). To integrate multilingual capabilities into existing LLMs, researchers have proposed a variety of methods to enhance multilingual adaptation. These approaches range from continual pretraining techniques (Ibrahim et al., 2024; Gupta et al., 2023) to initial training on extensive multilingual datasets (Scao et al., 2023; Chai et al., 2023) and then subsequent specialized fine-tuning on a target language (Yang et al., 2023; Han et al., 2022), and even adaptation through instruction tuning (Shaham et al., 2024; Kew et al., 2023; Gala et al., 2024). Critical aspects in multilingual adaptation remain on the availability of high-quality diverse multilingual corpus (Corrêa et al., 2024) and further the scope of vocabulary of the specific language.
Continual Pretraining. Static datasets are impractical for adapting to evolving real-world data, making continual learning essential (Ring, 1998; Thrun, 1998). Continual pretraining (Gururangan et al., 2020) allows models to incorporate new knowledge without retraining from scratch, a costly endeavor. As curated datasets like RedPajama (Together, 2023) and Dolma (Soldaini et al., 2024) become available, integrating them efficiently is crucial. This also enables the extension of models to new modalities, such as code (e.g., StableCode). Previous approaches focus on replay techniques, optimizing learning schedules (Ibrahim et al., 2024), soft masking (Ke et al., 2023), and forward/backward transfer (Yıldız et al., 2024).
@@ 7 Conclusion
In this work, we introduced Aurora-M, a multilingual model that extends the capabilities of code-focused LLMs while maintaining their original coding proficiency. We demonstrate that continual training from code to multilingual tasks is feasible, allowing the model to perform well across both domains. Adhering to the safety guidelines of the Biden-Harris US Executive Order on AI, Aurora-M promotes responsible AI development while pushing the boundaries of performance and utility. Our two-stage continual pretraining approach, combined with insights from cross-lingual red-teaming, highlights the adaptability and versatility of modern language models. Aurora-M serves as a valuable resource for both researchers and developers, fostering collaboration and transparency in the open-source AI community. Future work will explore continual pretraining on stronger base models with the same two-stage curriculum, focusing on safety for both LLMs and Multimodal-LLMs. We also aim to develop domain-specific expert models, enhancing task specialization and expanding model versatility.
Ethical Consideration
@@
Acknowledgments
This work was supported by the “R&D Hub Aimed at Ensuring Transparency and Reliability of Generative AI Models” project of the Ministry of Education, Culture, Sports, Science and Technology, and used resources of LUMI supercomputer under project_462000316.